If you added a new number to the number set, what value will the next number need to be for it to be thought of as an outlier?

data set: 12, 15, 19, 20, 22, 25, 28, 30
Answer #1

Answer)

12, 15, 19, 20, 22, 25, 28, 30, x

Let the new number be x

Now we have 9 data points

Median is the middlemost data point = 5th data point = 22

Q1 = middlemost data point of the numbers left to the median = 17

Q3 = middlemost data point of the numbers right to the median = 29

IQR = Q3 - Q1 = 29-17 = 12

Now anything above Q3 + (1.5*IQR) is an out lier

= 29 +(1.5*12)

= 47

So, the number can be any number above 47

Example : 48